OpenStudy (anonymous):

(2sqrt5 - sqrt3)(2sqrt3 + sqrt5)

OpenStudy (anonymous):

4+3sqrt15

OpenStudy (anonymous):

Thanks I need the steps how to get the answer though

OpenStudy (shamil98):

use foil

OpenStudy (anonymous):

Thats what Im doing at the moment. (2sqrt5)(2sqrt3) + (2sqrt5)(sqrt5) - (sqrt3)(2sqrt3) - (sqrt3)(sqrt5)

OpenStudy (anonymous):

The first part is 4sqrt15 right? But the second part Im not sure about...3sqrt5 or 2sqrt25???

OpenStudy (anonymous):

Nvm I figured it out thanks! 4sqrt15 + 2sqrt25 - 2sqrt9 - sqrt15 = 3sqrt15 +4!

OpenStudy (anonymous):

\[(2\sqrt{5}-\sqrt{3})(2\sqrt{3}+\sqrt{5})=(2\sqrt{5}(2\sqrt{3}+\sqrt{5})+(-\sqrt{3}(2\sqrt{3}+\sqrt{5})=(4*\sqrt{5}*\sqrt{3}+10)+(-2*\sqrt{3}*\sqrt{3}-\sqrt{3}*\sqrt{5})=3\sqrt{15}+4\]